one number is twice another number. If the larger is diminished by ten the result is two more than the smaller. Find the number.
Answer by oberobic(2304)

x is one number.
y is another.
We are told: y = 2x.
And we are told: y-10 = x+2
.
So, we have two definitions of y:
y = 2x
and
y-10 = x+2
y = x+12
Since the identity y=y is obvious, then we can say
2x = x+12
.
Subtracting x from both sides
x = 12
.
In our set up we defined: y = 2x, so y = 2*12 = 24
.
We can check by substituting these values into the other equation we defined:
y-10 = 14
Does this = 12+2?
Yes it does.
.
So, the two values are: 12 and 24.